OpenStudy (anonymous):

The point on the curve y=2x^3 - 33x^2 + 180x +1 whre the tangent line to it is parallel to the line joining (2,13) and (1,1)?

OpenStudy (experimentx):

find dy/dx .... equate it to slope ... solve for x

OpenStudy (anonymous):

find m, the slope of line through (2,13) and (1,1) y=2x^3 - 33x^2 + 180x +1 \[\frac{dy}{dx} = 6x^2 - 66x + 180 = m\] which is a quadratic. is this possible without calculus? @experimentX

OpenStudy (experimentx):

not sure ... if you like to try ... i'm ready to assist!!

OpenStudy (anonymous):

ahh unfortunately i must return to physics revision :( exam monday

OpenStudy (experimentx):

seems possible ... f(x1) - f(x2) = (x1 - x2)(some polynomial of x)

OpenStudy (anonymous):

i dint understand

OpenStudy (anonymous):

first i have to find the slope by y2-y1/x2-x1

OpenStudy (anonymous):

then dy/dx=0

OpenStudy (anonymous):

then wat to do?

OpenStudy (experimentx):

taking limit, we can replace (f(x1) - f(x2))/(x1 - x2) by m .. and x1 and x2 by x

OpenStudy (experimentx):

dy/dx = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

not possible without calc because "slope of the tannet line" is a calc concept

OpenStudy (anonymous):

*tangent

OpenStudy (anonymous):

ok @satellite73 can u show it plzz

OpenStudy (anonymous):

course i could be wrong

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

ah. i have seen methods using circles which work on certain curves though

OpenStudy (anonymous):

no i was agreeing with @eigenschmeigen take the derivative, set it equal to 12 and solve

OpenStudy (anonymous):

i'll try and find it

OpenStudy (anonymous):

i got x= 7 x =4

OpenStudy (anonymous):

is that correct

OpenStudy (experimentx):

find the corresponding y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thxxx @experimentX @satellite73 @eigenschmeigen

OpenStudy (experimentx):

lol ... yw, you find the tangent right??

OpenStudy (anonymous):

yeah..........